### v1

# Image is All You Need: Towards Efficient and Effective Large Language Model-Based Recommender Systems

Abstract.
Large Language Models (LLMs) have recently emerged as a powerful backbone for recommender systems. Existing LLM-based recommender systems take two different approaches for representing items in natural language, i.e., Attribute-based Representation and Description-based Representation. In this work, we aim to address the trade-off between efficiency and effectiveness that these two approaches encounter, when representing items consumed by users. Based on our interesting observation that there is a significant information overlap between images and descriptions associated with items, we propose a novel method, Image is all you need for LLM-based Recommender system (I-LLMRec). Our main idea is to leverage images as an alternative to lengthy textual descriptions for representing items, aiming at reducing token usage while preserving the rich semantic information of item descriptions. Through extensive experiments, we demonstrate that I-LLMRec outperforms existing methods in both efficiency and effectiveness by leveraging images. Moreover, a further appeal of I-LLMRec is its ability to reduce sensitivity to noise in descriptions, leading to more robust recommendations. Our code is available at https://github.com/rlqja1107/torch-I-LLMRec.
Recommender System, Large Language Models, Sequence modeling
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Large Language Models (LLMs) have recently demonstrated remarkable performance in various downstream tasks in the Natural Language Processing (NLP) domain, leveraging their strong semantic reasoning and world knowledge. These exceptional capabilities have stimulated researchers to explore their potential and broaden the application to diverse domains (Kim et al., 2024b; Chen et al., 2024; Kim et al., 2024a; Xie et al., 2024). In recommender systems, substantial efforts have also been made to utilize their rich semantic knowledge. Especially, recent studies explore the replacement of traditional collaborative filtering models (e.g., SASRec (Kang and McAuley, 2018) and LightGCN (He et al., 2020)) with LLMs as the backbone for recommendations (Lin et al., 2024; Bao et al., 2023; Zheng et al., 2024; Tan et al., 2024). To this end, they typically transform each item in a user’s interaction history from a traditional numerical ID into natural language (e.g., titles) and arrange them into sequences within the input prompt.
Figure 1. (a) Histogram of the input token length required to represent a user’s item interaction history across different item expression approaches. (b) Recommendation performance (Hit@5) and Inference Time (seconds/100 users) for different item representation approaches. We use the Amazon Sports dataset for analysis.
The key to LLM-based recommender systems lies in effectively representing items in natural language to capture user preferences based on LLM’s comprehensive understanding of user interaction history. In this regard, existing studies that represent items in natural language can be categorized into the following two main approaches: 1) Attribute-Based Representation approach is to combine simple attributes such as brand and category with the item title to represent the item (Bao et al., 2023; Li et al., 2023b; Tan et al., 2024; Lin et al., 2024). For example, TALLRec (Bao et al., 2023) represents items with their title for user preference modeling, while RECFORMER (Li et al., 2023b) enhances user preference understanding by adding brand and category attributes with attribute type-specific tokens in the prompt. Similarly, TransRec (Lin et al., 2024) incorporates attribute information alongside the title in the prompt to facilitate a multifaceted understanding of items. 2) Description-Based Representation approach is to use detailed item descriptions to preserve rich textual item semantics so that the LLM could capture user preference in a more fine-grained manner. For example, TRSR (Zheng et al., 2024) summarizes full descriptions using a large-scale LLM and provides them to a small-scale LLM for recommendation, thereby overcoming the limited item semantics provided by attributes only and addressing the input length constraints of LLMs (e.g., 1,024 tokens for GPT-2 (Radford et al., 2019)).
However, we argue that the two aforementioned approaches inherently face a trade-off between efficiency and effectiveness when representing items. Specifically, the Attribute-Based Representation approach is efficient since it requires fewer tokens to add item attributes than to add the entire item descriptions, which shortens the input length for the LLM [^1] Since LLMs are built on transformers, their computational complexity grows quadratically as the input token length increases.. However, it sacrifices the effectiveness of understanding user preferences due to the limited item semantics that can be contained in relatively short attributes. On the other hand, the Description-Based Representation approach is less efficient due to the longer input length required for detailed item descriptions, whereas the effectiveness of understanding user preferences is improved by utilizing richer textual semantics of items. In fact, as shown in Figure 1(a), the input length required for LLMs to represent a user’s item interaction history significantly varies depending on how the item is represented. That is, the Description-Based Representation approach (i.e., Summarized description [^2] To summarize the full description, we adopt an approach similar to TRSR (Zheng et al., 2024) and use a similar prompt. For more details, please refer to Appendix B. and Full description) demands much longer input token lengths than the Attribute-Based Representation approach (i.e., Attribute), resulting in the increased computational complexity.
To better understand this trade-off, in Figure 1(b), we extend our analysis beyond a simple comparison of input token length and examine the computational complexity of different item representation approaches (i.e., efficiency) along with their recommendation performance (i.e., effectiveness). Note that we use the same LLM backbone (i.e., Sheared LLaMA-2.7B (Xia et al., 2023)) and adopt the same recommendation protocol for fair comparisons (See Appendix C regarding the details of recommendation protocol). We observe that the Attribute-based Representation approach (blue circle) achieves the inference time more than 2.5 times faster than the Description-based Representation approach (green and black). However, this comes at the expense of a performance reduction of over 13% due to the lack of rich item semantics contained in attributes compared with descriptions. Furthermore, summarizing full descriptions reduces inference time but compromises performance due to the information loss during the summarization process. In conclusion, the trade-off can be summarized as: richer (i.e., longer) item representation provided to the LLM improves effectiveness but decreases efficiency. However, this trade-off is unavoidable by nature if items are to be represented with natural language.
Figure 2. Cosine similarity between item image-description pairs in Amazon Sport and Art datasets and image-caption pairs in the COCO dataset using CLIP (Radford et al., 2021).
In this paper, we focus on addressing this trade-off based on our interesting observation that there is a significant information overlap between images and descriptions associated with items. Specifically, when we measure the similarity between item image-description pairs in a real-world e-commerce dataset (i.e., Amazon Sport and Art (Ni et al., 2019)) using a vision-language model (i.e., CLIP (Radford et al., 2021)) in which image and language spaces are jointly embedded, we found that the similarity is surprisingly high. Specifically, Figure 2 shows that the average similarity for the Amazon Sport and Art datasets is around 0.31. To better understand this similarity level, we conducted the same experiment with the COCO caption dataset (Lin et al., 2014), a well-curated image-caption pair dataset widely used in vision-language research. Specifically, we measured similarities for both positive (i.e., image-caption) and negative (i.e., image-randomly sampled unrelated caption) pairs. Given that the negative pairs have an average similarity of only 0.07, the positive pairs’ average similarity of around 0.26 can be considered an indicator of high similarity. Furthermore, the similarity value of 0.31 between item image-description pairs, which is even higher, highlights a significant information overlap between images and their associated descriptions [^3] Such an observation was consistently observed across other real-world datasets as well as another CLIP variant model (See Appendix D)..
Building on this observation, we propose a novel method, Image is all you need for LLM-based Recommender system (I-LLMRec), which addresses both efficiency and effectiveness of existing LLM-based recommender systems by leveraging images as an alternative to lengthy textual descriptions for the item expression, aiming at reducing token usage while preserving the rich semantic information of item descriptions [^4] Some text-specific information that is missing in an image may be lost, but we find that in Section 4.2, this information has surprisingly little impact on recommendations.. The main technical challenge lies in the misalignment between the item image space and the language space. To this end, we adopt a learnable adaptor for visual features followed by a technique to bridge the gap between the two spaces (i.e., Image-LLM Alignment mode (ILA)), which facilitates the training of the adaptor with carefully crafted prompts tailored to the recommendation context. This ensures the LLM to capture rich item semantics through images with only a few tokens, thereby effectively and efficiently capturing user preferences.
Through extensive experiments on various real-world datasets, we demonstrate that our proposed method using images instead of item descriptions is superior in both effectiveness and efficiency. Specifically, I-LLMRec improves inference speed by approximately 2.93 times compared to the Description-based Representation approach, while achieving a 22% performance improvement over the Attribute-based Representation approach (Figure 1(b)). As a further appeal, I-LLMRec facilitates robust recommendations by mitigating sensitivity to noise associated with item descriptions as it relies on images. Our main contributions can be summarized as follows:
  • We identify a trade-off between efficiency and effectiveness when expressing items in natural language.
  • Based on the observation that there is a significant overlap between item image-description pairs, we propose a novel method, called I-LLMRec, which utilizes images instead of textual descriptions to efficiently and effectively capture user preferences.
  • Our extensive experiments demonstrate that I-LLMRec outperforms the various natural language-based representation approaches in both effectiveness and efficiency.
2. Preliminary
In this section, we first introduce the task formulation and then examine the complexity of different item expression approaches.
Task Formulation. Throughout this paper, we mainly focus on sequential recommendation, as it closely aligns with real-world scenarios (Tian et al., 2022; Xie et al., 2022). Let $\mathcal{U}$ and $\mathcal{I}$ denote the set of users and items, respectively. A user $u\in\mathcal{U}$ has the historical item interaction sequence denoted as $\mathcal{S}_{u}$=[$i_{1},i_{2},...,i_{k},...,i_{|\mathcal{S}_{u}|}$], where $i_{k}$ denotes the $k$-th interacted item and $|\mathcal{S}_{u}|$ is the number of items in the interaction sequence. The goal of this task is, for each user $u$, to predict the next item $i_{|\mathcal{S}_{u}|+1}$ to be consumed by the user based on the user interaction history $\mathcal{S}_{u}$.
Representing Item Semantics. With the advancement of LLMs for recommender systems, there is a growing need to help them understand the semantics of items to capture user preferences. In this regard, we categorize the multiple types of information to recommend an item $i$’s semantics into [$\mathbf{I}_{i}$, $\mathbf{D}_{i}$, $\mathbf{A}_{i}$], where $\mathbf{I}_{i}$, $\mathbf{D}_{i}$, and $\mathbf{A}_{i}$ are an item image, textual description [^5] To avoid confusion, we will henceforth refer to the description as summarized description instead of full description unless stated otherwise., and attribute (e.g., brand and category), respectively.
Comparison of complexity. Let $f(\cdot)$ denote the transformation function for the LLM to convert the representation of item semantics into input tokens, and let $|f(\cdot)|$ denote the number of input tokens. Specifically, for $\mathbf{D}_{i}$ and $\mathbf{A}_{i}$ that are in natural language, $f$ refers to tokenizing the natural language and converting them into the word embeddings. Meanwhile, for an image $\mathbf{I}_{i}$, $f$ involves extracting the visual features using a pretrained vision encoder (e.g., CLIP-ViT (Radford et al., 2021)), followed by an adaptor to resize their feature dimensions for compatibility with the LLM (Liu et al., 2024b; Li et al., 2023a). The complexity of LLM operations for each representation is $O((|f(\mathbf{I}_{i})||\mathcal{S}_{u}|)^{2}d)$, $O((|f(\mathbf{D}_{i})||\mathcal{S}_{u}|)^{2}d)$, and $O((|f(\mathbf{A}_{i})||\mathcal{S}_{u}|)^{2}d)$, where $|f(\mathbf{D}_{i})|>>|f(\mathbf{A}_{i})|$. This is derived based on the complexity of Transformer, i.e., $O(n^{2}d)$, where $n$ is the input token length and $d$ is the feature dimension of the LLM. It reveals that the complexity of the description-based approach is not merely being incurred with a single item; rather, it increases with the user sequence length $|\mathcal{S}_{u}|$, and its complexity even grows quadratically, making this approach impractical. Therefore, in this work, we propose to lower the complexity of expressing an item by leveraging its associated image using only a single token (i.e., $|f(\mathbf{D}_{i})|\gg|f(\mathbf{A}_{i})|>|f(\mathbf{I}_{i})|=1$ ) [^6] In average, $|f(\mathbf{D}_{i})|$ and $|f(\mathbf{A}_{i})|$ are approximately 160 and 10 tokens, respectively. while preserving the rich semantics of the item description. This is based on our observation that there is a significant information overlap between the item image-description pairs (Figure 2).
3. Proposed Method: I-LLMRec
In this section, we provide a detailed explanation of I-LLMRec. After introducing the initial recommendation-related task prompt, we begin by encoding the images of items consumed by users, and mapping them to an LLM through a learnable adaptor (Section 3.1). However, since these mapped visual features are not inherently aligned with the language space, we propose the Image-LLM Alignment (ILA) module that trains the adaptor with carefully crafted prompts tailored for the recommendation context, resulting in effectively bridging the alignment gap (Section 3.2). Meanwhile, we propose the Image-based REtrieval (IRE) module, a training approach that enables the LLM to directly perform recommendations from the item corpus based on the visual features (Section 3.3). Finally, we describe the overall training objectives and inference for recommendation (Section 3.4 and 3.5). The overall framework of I-LLMRec is shown in Figure 3.
3.1. Mapping Item Images to LLM
Recall that, as shown in Figure 2, the item image contains rich item semantics of the lengthy description. Therefore, in this section, our goal is to map items consumed by users to an LLM using only a few tokens, allowing us to capture user preference efficiently and effectively. Specifically, given the interaction history of user $u$, i.e., $[i_{i},i_{2},...,i_{|\mathcal{S}_{u}|-1}]$, and a target item $i_{|\mathcal{S}_{u}|}$, we convert the interaction history into a sequence of item images $[\mathbf{I}_{i_{1}},\mathbf{I}_{i_{2}},...,\mathbf{I}_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times H\times W\times 3}$, where $H$ and $W$ are the image height and width, respectively. Then, we adopt a pretrained visual encoder $V$ of a vision-language model to extract the visual features of each item, defined as $v_{i}=V(\mathbf{I}_{i})\in\mathbb{R}^{d_{v}}$. This process makes the sequence of visual features $[v_{i_{1}},v_{i_{2}},...,v_{i_{|\mathcal{S}_{u}|-1}}]\in\mathbb{R}^{(|\mathcal{S}_{u}|-1)\times d_{v}}$, where $d_{v}$ is the dimension of the visual feature.
3.1.1. Adaptor network for visual features.
However, since the visual feature dimension $d_{v}$ is different from that of the LLM feature $d$, and their spaces are inherently misaligned, we introduce an adaptor network $M:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d}$ to make visual features compatible with the LLM’s feature dimension, allowing $v_{i}$ to be interactive with word embeddings within the LLM’s input layer. More precisely, we process each item’s visual feature $v_{i}$ through the adaptor $M$, transforming it into a sequence of visual features that match the dimensionality of the LLM. Formally, we the sequence of mapped visual features is defined as:
(1) $\mathcal{\bar{S}}_{u}^{\mathbf{I}}=[\bar{v}_{i_{1}},\bar{v}_{i_{2}},...,\bar{v}_{i_{|\mathcal{S}_{u}|-1}}],\bar{v}_{i}=M(v_{i})\in\mathbb{R}^{d}.$
Figure 3. Overall framework of I-LLMRec. User-interacted item images are mapped into the LLM through an adaptor, which bridges the image and language spaces. To ensure alignment between two spaces, the adaptor is optimized via the ILA module. Furthermore, the recommendation process is formulated as a retrieval task via the IRE module.
3.1.2. Designing a prompt to represent a user’s item interaction history.
To represent a user’s item interaction history with visual features, we carefully design a prompt, enabling the LLM to comprehensively understand the item semantics. Specifically, for each item $i$ in a user’s item interaction history, we represent it in the prompt as follows:
[…]
3.2. Image-LLM Alignment (ILA) Module
To make the LLM effectively capture user preferences from item images, it is crucial to align visual features with the language space in a meaningful way, requiring adequate supervision of the adaptor network $M$. To achieve this, we optimize the adaptor network $M$ by predicting the next item properties based on the user interaction within a structured prompt. This approach ensures that the LLM learns to associate visual features with relevant textual information, thereby improving its ability to understand user preferences in the context of recommendation.
Specifically, we craft a prompt in an "input"-"target output" format. Here, the input consists of the user interaction prompt, followed by a question regarding the next item (e.g., What brand is this user likely to consume for the next item?). The target output corresponds to the answer to that question (e.g., The brand likely to be consumed is ‘WOLT’). To guide the LLM in considering various properties of the next item, we incorporate multiple properties of the next item, including brand, category, title, and description. For each property, we design five different question templates (refer to Appendix I for complete templates). During each training step, we randomly select one of 20 possible question templates (4 properties × 5 templates) and train $M$ to generate the corresponding output. The training objective is formulated as follows:
[…]
where $\theta$ is the parameter of the LLM that is frozen, $y$ is the target output, $y_{k}$ is the $k$-th token of $y$, and $x$ is the input.
3.2.1. Discussion regarding LLM fine-tuning
A further benefit of I-LLMRec is that LLM fine-tuning is not mandatory, unlike the existing LLM-based recommender systems (Lin et al., 2024; Bao et al., 2023; Tan et al., 2024; Zheng et al., 2024; Liu et al., 2024a) that generally rely on costly fine-tuning of the LLM. Instead, we harness the LLM’s intrinsic parametric knowledge to understand the user preferences. Even though I-LLMRec could be potentially improved by fine-tuning the LLM (See Appendix E), we opt to keep the LLM frozen to preserve the LLM’s intrinsic knowledge and improve the training efficiency.
3.3. Image-based REtrieval (IRE) Module
We found that existing studies on LLM-based recommender systems often face challenges in providing reliable and efficient recommendations. This is because they typically recommend items by predicting the next item title (Kim et al., 2024a; Lin et al., 2024; Tan et al., 2024) or item tokens allocated as out-of-vocabulary (Wei et al., 2024; Li et al., 2023b; Zhu et al., 2024). Regarding reliable recommendation, the title prediction approach cannot guarantee that the recommended items exist within the item corpus (i.e., lack of item grounding), which sometimes may lead to recommending non-existent items. Regarding efficient recommendation, title prediction relies on computationally intensive beam search to recommend multiple items, while the item token prediction approach requires an oversized LLM, as expanding the item pool necessitates extending the LLM vocabulary set, which hinders scalability. To overcome these challenges, we propose the Image-based REtrieval (IRE) module, which formulates the recommendation task as a retrieval task by leveraging the images to directly retrieve relevant items from the item pool. In the following, we describe how to obtain an LLM-guided user representation for retrieval, and define a training objective that retrieves relevant items using the visual features.
3.3.1. LLM-guided user representation
[…]
3.3.2. Training objective for retrieving relevant items
[…]
(4) $o_{u}^{\textsf{Img}}=F^{\textsf{Img}}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{\textsf{Img}}=F^{\textsf{Img}}_{i}(v_{i_{|\mathcal{S}_{u}|}})$
where $F^{\textsf{Img}}_{u}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{s}}$, $F^{\textsf{Img}}_{i}:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{s}}$ are the projectors, and $d_{s}$ is the feature dimension of shared recommendation space. $o_{u}^{\textsf{Img}}$ and $o_{i}^{\textsf{Img}}$ denote the projected representations of user $u$ and item $i$ in terms of visual features, respectively. To optimize the retrieval task, we use the binary cross-entropy loss as the training objective:
(5) $\begin{split}\mathcal{L}_{\text{IRE}}^{\textsf{Img}}&=-\sum_{u\in\mathcal{U}}\left[log(\sigma(r_{u,i}^{\textsf{Img}}))+log(1-\sigma(r_{u,i^{-}}^{\textsf{Img}}))\right]\end{split}$
[…]
3.3.3. Extension to multiple feature types.
Our retrieval-based recommendation approach can seamlessly incorporate multiple item feature types (e.g., ID-based embeddings and textual features) alongside visual features. Specifically, for a given item feature type denoted as *, we can simply integrate it by introducing two additional projectors, $F^{*}_{u}$ and $F^{*}_{i}$:
(6) $o_{u}^{*}=F^{*}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{*}=F^{*}_{i}(\mathbf{IF}^{*})$
where $\mathbf{IF}^{*}$ represents the item feature corresponding to feature type * (e.g., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$ if * is the image type Img), and $o_{u}^{*}$ and $o_{i}^{*}$ are the projected representations of user $u$ and item $i$ in terms of * feature type, respectively. Using $o_{u}^{*}$ and $o_{i}^{*}$, we can then compute the binary cross-entropy loss $\mathcal{L}_{\text{IRE}}^{*}$ following Equation 5.
Among various item feature types, we extract the ID-based item embeddings from a pretrained collaborative filtering (CF) model (i.e., SASRec (Kang and McAuley, 2018)), thanks to its effectiveness in general recommendation tasks, especially for warm items (Kim et al., 2024a) [^7] In Figure 7, we show that the utilization of ID-based item embeddings leads to more recommendations of warm items.. Specifically, we incorporate ID-based item embeddings (i.e., $c_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{CF}}$) in addition to the visual features (i.e., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$). Furthermore, as item descriptions are readily available in real-world scenarios (Zheng et al., 2024), we can easily incorporate textual features $t$ extracted from full descriptions to further improve overall performance (i.e., $t_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Text}}$). The analysis of extension to various features is provided in Section 4.4.
3.4. Model Training
[…]
(7) $\begin{split}\mathcal{L}_{final}=\mathcal{L}_{\text{ILA}}+\mathcal{L}_{\text{IRE}}^{\textsf{Img}}+\mathcal{L}_{\text{IRE}}^{\textsf{CF}}+\mathcal{L}_{\text{IRE}}^{\textsf{Text}}\end{split}$
3.5. Inference
For retrieval-based inference, we compute the affinity scores $r_{u,i}^{\textsf{Img}}$, $r_{u,i}^{\textsf{CF}}$, and $r_{u,i}^{\textsf{Text}}$ using the scoring function $\mathcal{T}$, in terms of visual, CF, and textual features, respectively. Specifically, the top-$k$ relevant items for $i_{|\mathcal{S}_{u}|+1}$ is computed as follows:
(8) $rec_{u}^{k}=\textsf{Top-k}(r_{u,i}^{\textsf{Img}}+r_{u,i}^{\textsf{CF}}+r_{u,i}^{\textsf{Text}}),\quad\forall i\in\mathcal{I}$
where $rec_{u}^{k}$ is the set of recommended $k$ items for user $u$, and Top-k is a function that extracts items with the highest top-$k$ scores. Note that while the affinity scores across different features can be aggregated through various approaches, such as weight summation or product, we opt for a simpler summation.
Table 1. Performance Comparison. The best performance is denoted in bold text, while the second-best among baselines is underlined. $\mathbf{A}$: Attributed-based Representation, $\mathbf{CF}$: CF-based Representation (i.e., the CF item embedding is projected into the LLM), $\mathbf{D}$: Description-based Representation, $\mathbf{I}$: Image-based Representation.
Dataset Metric Collaborative Filtering LLM Image-aware LLM
GRU4Rec VBPR BERT4Rec SASRec TALLRec ($\mathbf{A}$) A-LLMRec ($\mathbf{CF}$) TRSR ($\mathbf{D}$) UniMP ($\mathbf{I}$) I-LLMRec ($\mathbf{I}$)+$\mathbf{D}$ I-LLMRec ($\mathbf{I}$)
Sport NDCG@5 0.2106 0.2369 0.2389 0.3129 0.2938 0.3352 0.3375 0.3364 0.3637 0.3711
Hit@5 0.2820 0.3097 0.2993 0.3841 0.3801 0.4070 0.4302 0.4030 0.4554 0.4570
NDCG@10 0.2413 0.2694 0.2670 0.3514 0.3323 0.3683 0.3765 0.3629 0.4003 0.4071
Hit@10 0.3775 0.4108 0.3866 0.4760 0.4997 0.5096 0.5515 0.4853 0.5694 0.5689
Grocery NDCG@5 0.2673 0.2321 0.2995 0.3753 0.3477 0.3860 0.3802 0.3710 0.3908 0.3956
Hit@5 0.3609 0.3171 0.3834 0.4684 0.4589 0.4823 0.4917 0.4506 0.5037 0.5069
NDCG@10 0.3033 0.263 0.3317 0.4096 0.3874 0.4195 0.4184 0.4011 0.4288 0.4332
Hit@10 0.4726 0.4232 0.4835 0.5746 0.5815 0.5864 0.6099 0.5439 0.6211 0.6232
Art NDCG@5 0.3119 0.3710 0.3626 0.4561 0.4572 0.4652 0.4758 0.4565 0.4796 0.4839
Hit@5 0.4044 0.4656 0.4455 0.5374 0.5663 0.5681 0.5841 0.5315 0.5902 0.5883
NDCG@10 0.3481 0.4062 0.3955 0.4860 0.4944 0.4981 0.5100 0.4829 0.5160 0.5191
Hit@10 0.5203 0.5745 0.5477 0.6301 0.6813 0.6877 0.6896 0.6131 0.6981 0.6974
Phone NDCG@5 0.2023 0.1898 0.2319 0.3299 0.3721 0.3403 0.3886 0.3388 0.3892 0.3900
Hit@5 0.2877 0.2688 0.3184 0.4366 0.4986 0.4502 0.5148 0.4427 0.5176 0.5156
NDCG@10 0.2353 0.2222 0.2664 0.3658 0.4148 0.3811 0.4292 0.3757 0.4309 0.4320
Hit@10 0.3952 0.3698 0.4255 0.5478 0.6305 0.5761 0.6401 0.5569 0.6463 0.6448
4. Experiments
In this section, we conduct experiments to explore the following research questions:
  • RQ1: How does I-LLMRec perform compared to the CF and LLM-based recommender models?
  • RQ2: How well does I-LLMRec offer strengths (i.e., efficiency, effectiveness, and robustness) by utilizing images rather than lengthy descriptions?
  • RQ3: How does each module (i.e., ILA and IRE) contribute to the effectiveness of I-LLMRec?
  • RQ4: How can we handle when item images are missing?
4.1. Experimental Setting
Datasets. For evaluation, we use four categories from the Amazon dataset (Ni et al., 2019): Sports, Grocery, Art, and Phone. These datasets contain the item titles, attributes (e.g., brand and category), detailed textual descriptions, and images representing items. Following a prior study (Wei et al., 2024), we filter out users and items with fewer than 5 interactions to ensure data quality. We summarize the data statistics in the Appendix F.
Evaluation Protocol. Following the leave-one-out protocol (Kim et al., 2023; Kang and McAuley, 2018), we use each user’s most recent interaction as the test set, the second most recent interaction as the validation set, and the remaining interactions as the training set. For evaluation metrics, we utilize Hit Ratio (Hit@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$) with $k=5,10$. Hit@$k$ measures whether the ground-truth item appears in the recommended list (i.e., $rec_{u}^{k}$) while NDCG@$k$ evaluates its ranking position. To reduce computational complexity during evaluation for each user, we randomly select 100 negative items that the user has not interacted with, add the ground-truth test item, and compute the metrics based on this set.
Baselines. We compare I-LLMRec with the following baselines:
Collaborative Filtering (CF) models.
  • GRU4Rec (Hidasi, 2015) employs the RNN network to capture sequential user interactions.
  • VBPR (He and McAuley, 2016b) enhances BPR (Rendle et al., 2012) by incorporating image features to improve personalized ranking.
  • BERT4Rec (Sun et al., 2019) applies bidirectional transformers with masked item prediction to capture complex user preference.
  • SASRec uses self-attention mechanisms to capture long-term user preference.
LLM-based models.
  • TALLRec (Bao et al., 2023) converts numerical IDs into titles while intentionally adding attributes (i.e., brand and category), regarding it as a representative model of the Attribute-based Representation approach.
  • A-LLMRec (Kim et al., 2024a) express items by projecting the CF item embeddings into the LLM along with title, enabling effective recommendation of warm items.
  • TRSR (Zheng et al., 2024) is a representative work of the Description-based Representation approach, which summarizes full item descriptions to effectively express items while leveraging rich textual semantics.
Image-aware LLM-based models.
  • UniMP (Wei et al., 2024) utilizes image features and attributes to represent items, and recommends them through image tokens assigned as out-of-vocabulary.
  • I-LLMRec+$\mathbf{D}$ (equiv. to I-LLMRec+TRSR) [^8] For each item in the user interaction, we append the description as ”Description: $\mathbf{D}_{i}$” after $\mathbf{P}_{i}$ within the prompt, i.e., $\mathbf{P}_{i}=\{\textrm{Title}:\textsf{ITEM\_TITLE},\textrm{Visual}\;\textrm{Representation}:[\textsf{VISUAL}],\textrm{Description}:\textbf{D}_{i}$ is a variant of I-LLMRec that represent items using both image features and text description.
Note that the main difference between LLM-based models and I-LLMRec lies in how item semantics are represented (i.e., attributes, CF, descriptions, or images) in that both share the same LLM backbone and recommendation protocol.
4.1.1. Implementation Details
We employed Sheared LLaMA 2.7B (Xia et al., 2023) as the LLM backbone. For the visual encoder $V$, we adopt SigLIP (Zhai et al., 2023), a variant of CLIP (Radford et al., 2021) with a sigmoid loss function, to extract visual features. Regarding experiments with different LLMs and visual encoders, please refer to Appendix G. We use a 2-layer MLP with ReLU activation function for the adaptor $M$. In IRE module, we employ SBERT (Reimers, 2019) as the textual encoder. To ensure consistency across the LLM backbone, we use the same Sheared LLaMA 2.7B for TALLRec, A-LLMRec, and TRSR. Since these models are limited to recommending items only from a narrower item pool (i.e., 1 or 20), we apply the same IRE module for recommendation, which incorporates image, CF, and textual features, allowing them to recommend items from a larger pool. This approach is acceptable since our main goal is to evaluate how effectively the LLM captures user preferences through item expression. For TRSR, we use GPT-4o (OpenAI, 2024) to summarize full item descriptions. For UniMP, we follow its original implementation, which utilizes OpenFlamingo-4B-Instruct (Awadalla et al., 2023), a pre-trained multimodal LLM. Regarding training configurations, we set the learning rate to 0.001 and the batch size to 32. All models are trained on an NVIDIA GeForce A6000 48GB GPU.
4.2. Performance Comparison (RQ1)
Table 1 shows the performance comparison between I-LLMRec and baselines across four datasets. The key observations are as follows: 1) LLM-based models generally outperform traditional CF models, highlighting the effectiveness of leveraging an LLM backbone for the recommender systems. 2) A-LLMRec and TRSR generally outperform TALLRec. It suggests that beyond the attributes, incorporating CF item embeddings or item descriptions further enhances the LLM’s ability to capture user preferences. However, we observe that A-LLMRec performs worse on the Phone dataset, and TRSR generally outperforms A-LLMRec, indicating that item descriptions are more effective than ID-based item embeddings in expressing item semantics. 3) I-LLMRec outperforms UniMP, an image-aware LLM-based model. Given that UniMP employs a multimodal LLM where images and texts are pre-aligned, it indicates the need for a more specialized alignment strategy tailored to the recommendation tasks, demonstrating the effectiveness of I-LLMRec. 4) Considering that TRSR includes attributes during summarization (See Appendix B), adding description information in TALLRec, which is equivalent to TRSR, improved performance, whereas descriptions in addition to images (i.e., I-LLMRec+$\mathbf{D}$) did not improve the performance of I-LLMRec. This suggests that while descriptions convey useful information, they are inherently limited by what is already present in images due to the significant information overlap between image and description pairs. Even if descriptions may include some text-specific information that images do not contain, the fact that performance remains competitive implies that such information has surprisingly little impact on the recommendations. This supports the idea that item images are sufficient to capture user preferences for recommendation.
4.3. Model Analysis (RQ2)
In this section, we analyze how I-LLMRec, which represents items using images, performs in recommendations compared to the natural language-based approach (TRSR, I-LLMRec+$\mathbf{D}$, and TALLRec) in terms of efficiency, effectiveness, and robustness.
Figure 4. Inference time (seconds/100 users) over the length of users’ item interaction sequence (i.e., $|\mathcal{S}_{u}|$).
4.3.1. Analysis on Efficiency.
To study the efficiency of I-LLMRec, we evaluate the inference time across different sizes of user’s item interaction sequences (i.e., $|\mathcal{S}_{u}|$). Specifically, we divided the users into groups according to $|\mathcal{S}_{u}|$, randomly selected 100 users per group, and measured their total inference time. As shown in Figure 4, we observe that TRSR and I-LLMRec+$\mathbf{D}$, which rely on lengthy descriptions, show a sharp increase in inference time as the length of textual descriptions within the prompt increases proportional to $|\mathcal{S}_{u}|$. On the other hand, I-LLMRec maintains consistently low inference time regardless of $|\mathcal{S}_{u}|$. This efficiency stems from the fact that even as $|\mathcal{S}_{u}|$ increases, only a minimal number of image tokens are added, keeping computational costs low. Furthermore, we observe that TALLRec, which is an attribute-based representation approach, shows a gradual increase in inference time and eventually surpasses I-LLMRec when $|\mathcal{S}_{u}|\geq 10$. We attribute this to the fact that processing natural language attributes scales in complexity faster than image processing required for I-LLMRec. In summary, I-LLMRec is more efficient than TRSR across all $\mathcal{S}_{u}$ and even more efficient than TALLRec when user interactions are high, demonstrating superior computational efficiency.
Figure 5. Performance of various LLM context window sizes.
4.3.2. Analysis on Effectiveness.
To gain a deeper understanding of the effectiveness of I-LLMRec, we examine how well it performs under a limited budget compared to the natural language-based approach. Specifically, we consider the LLM’s context window size (i.e., maximum input token length) as the budget, and vary it from 4,096 to 256 tokens. This allows us to analyze the ability of models in effectively capturing the user preferences given a limited budget. As shown in Figure 5, we observe that TRSR, I-LLMRec+$\mathbf{D}$, and I-LLMRec perform consistently when the context window size is over 1,024 since they can handle most interactions. However, as the context window size drops below 512, the performance of TRSR and I-LLMRec+TRSR declines sharply. This is because lengthy descriptions make it difficult for the LLM to fully process interactions with a limited context window size, thereby making it more challenging to understand user preferences. On the other hand, I-LLMRec maintains stable performance even at the context window size of 256 by representing item semantics using minimal image tokens. This enables the LLM to process full user interactions within the given budget and effectively capture user preferences. While TALLRec also remains stable at 256 due to its low token usage, it exhibits inferior performance because its attributes are not expressive enough. Overall, I-LLMRec effectively overcomes the context window size constraints by leveraging item images to preserve rich item semantics, demonstrating its effectiveness and practicality.
Figure 6. Performance of natural language-based approaches (TALLRec, TRSR and Full Description) and I-LLMRec on four datasets (Sports, Grocery, Phone and Art).
4.3.3. Analysis on Robustness
To evaluate the robustness of representing items using images instead of natural language as done in I-LLMRec, we compare its performance with various natural language-based approaches—TALLRec (Attribute), TRSR (Summarized description), and Full description—across multiple datasets. Figure 6 presents two key findings regarding the drawback of natural language-based approaches. 1) Information loss from summarization (Figure 6(a)): Full description outperforms TRSR, indicating that summarizing descriptions introduces information loss, where essential semantic details may be omitted, leading to lower performance. 2) Presence of Noise in Full Description (Figure 6(b)): On some datasets (i.e., Phone and Art), Full description rather performs worse than TRSR that relies on a summarized description. This indicates that the lengthy full description may contain noise and that the noise could be partially removed in the summarized version of the full description [^9] Indeed, we found that full descriptions often contain extraneous content, such as HTML tags introduced during data crawling.. Especially in Phone dataset, Full Description even underperforms TALLRec that relies only on simple attributes. Such different behaviors of natural language-based approaches across different datasets demonstrates that these approaches are sensitive to the given text, and thus impractical in reality. In contrast, I-LLMRec, which captures item semantics through images rather than text, consistently delivers strong performance across all datasets, demonstrating its robustness in handling variations in item descriptions.
Table 2. Ablation studies of I-LLMRec.
Row ILA IRE Sport Art
Image CF Text Hit@5 NDCG@5 Hit@5 NDCG@5
(a) 0.3953 0.3043 0.5040 0.3915
(b) 0.4316 0.3403 0.5564 0.4447
(c) 0.4075 0.3256 0.5502 0.4517
(d) 0.4491 0.3630 0.5795 0.4769
(e) 0.4570 0.3711 0.5883 0.4839
Figure 7. Performance across item groups, where higher IDs indicate warmer item groups.
4.4. Ablation Studies (RQ3)
In Table 2, we conduct ablation studies to understand the effectiveness of each component in I-LLMRec. The variant without any additional components is trained solely using Equation 5 and recommends items based only on image features (row (a)). We make the following observations: 1) Effect of ILA: We observe that equipment with the ILA module significantly improves performance (row (a) vs. row (b)). This indicates that the ILA module effectively aligns visual features with the language space, enhancing the LLM’s ability to understand user preferences. 2) Effect of extending to multiple feature types in IRE: We observe that incorporating CF features along with image features [^10] The inclusion of specific feature types is used for both the model training in Equation 7 and during inference in Equation 8. yields better performance than using either the image or CF feature alone (row (b), (c) vs. (d)). Moreover, integrating all three feature types (i.e., Image, CF, and Text) further improves performance (row (d) vs. (e)). This demonstrates that extending multiple feature types enhances recommendation effectiveness beyond image features alone.
Effectiveness of Feature types across cold/warm items. To further analyze the impact of different feature types, we evaluate their performance on cold and warm item recommendations. Specifically, we follow (Liu et al., 2023) by sorting the items in ascending order based on the frequency of interaction and dividing them into five equally sized groups, where a higher group ID indicates a warmer item group, Then, we compare the recommendation performance across these groups. As shown in Figure 7, we observe that using image features (blue bar) generally outperforms CF features (orange bar) for cold items (lower ID groups). However, as the item group transitions from cold to warm, CF’s effectiveness improves, eventually surpassing Image in Item Group 5 (Warm). This suggests that CF is particularly effective for recommending warm items. With these insights, combining Image and CF (green bar) helps mitigate CF’s weakness in recommending cold items, resulting in higher performance than CF alone in Item Group 1-2. At the same time, the CF features compensate for weakness of the image features in recommending warm items, leading to better performance than image alone in Item Group 4-5. Furthermore, adding textual features (red bar) further boosts performance across all Item Groups, exceeding the performance of Image+CF. This analysis highlights how different feature types complement each other to enhance recommendation.
Table 3. Performance when item images are missing.
Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
TALLRec 0.3801 0.2938 0.5663 0.4572
TRSR 0.4302 0.3375 0.5841 0.4758
I-LLMRec w/ missing images 0.4451 0.3589 0.5805 0.4789
4.5. Missing Image Scenario (RQ4)
In this section, we explore how to handle scenarios where item images are missing and evaluate how effectively we can manage them. Specifically, we randomly remove 50% of the item images from the entire item pool. As a result, these missing images cannot be used for item semantics representation (Equation 2) and retrieval-based recommendation (Equation 5 and Equation 8). To compensate for the missing images, we substitute them with readily available textual descriptions. Similarly, for retrieval-based recommendations, when item images are missing, we extract textual features derived from descriptions using SigLIP (Zhai et al., 2023) text encoder rather than extracting visual features from images using SigLIP visual encoder. As shown in Table 3, I-LLMRec trained with half of the items missing images outperforms the baselines (TALLRec and TRSR), even though the baselines were trained without any missing item images. This demonstrate the effectiveness of I-LLMRec in handling the missing image scenario, which shows the practicality of I-LLMRec in reality.
5. Related Work
Sequential Recommendation. Our setup closely aligns with the sequential recommendation, where a user’s item interaction history is listed as a sequence in chronological order, and the goal is to predict the user’s next interaction item. Early studies process user sequences via the Markov chain for sequential recommendation (He and McAuley, 2016a; Mahmood and Ricci, 2007; Wang et al., 2015). With the advancement of deep learning, RNNs were used for sequence modeling to capture user preference (Hidasi, 2015; Li et al., 2017), while studies using CNN treat previous interacted items’ embedding matrices as images, enabling the convolutional operation to consider user sequences (Tang and Wang, 2018; Yuan et al., 2019; Kim et al., 2016). Recently, with the emergence of Transformer (Vaswani, 2017), the self-attention mechanism has been applied to sequential recommendations (Kang and McAuley, 2018; Sun et al., 2019; Xu et al., 2021). More recently, the focus has shifted towards leveraging rich side information associated with items (e.g., images or text) in the sequential recommendation. Specifically, TempRec (Wu et al., 2022) encodes textual information of items to strengthen item embeddings, while MMMLP (Liang et al., 2023) fuses visual and textual features in the user sequence to capture fine-grained user preferences.
LLM-based Recommendation. Due to the power of textual understanding and world knowledge of LLMs, a myriad of studies have spurred the development of LLMs for recommender systems (Bao et al., 2023; Wang and Lim, 2023; Kim et al., 2024a; Zhang et al., 2023; Liu et al., 2024a). To adapt LLMs for recommendation tasks, previous studies represent the interaction history as a sequence of item titles. Additionally, some studies enhance recommendation performance by incorporating item attributes, providing richer item semantics to the LLMs (Attribute-based Representation). Specifically, TALLRec (Bao et al., 2023) converts item IDs into their titles and captures user preference via LoRA (Hu et al., 2021) fine-tuning. TransRec (Lin et al., 2024) improves the user preference understanding by including item titles and attributes in prompts. IDGenRec (Tan et al., 2024) incorporates item titles generated by an independent LLM for unique and meaningful semantics. More recently, to feed the richer item semantics to LLMs, a description of items have been utilized (Description-based Representation). Specifically, TRSR (Zheng et al., 2024) utilizes a large LLM (i.e., LLaMA-30B-instruct (Touvron et al., 2023a)) to summarize item descriptions, and feeds them to a smaller LLM (i.e., LLaMA-2 7B (Touvron et al., 2023b)) to capture the condensed meaningful information, while ONCE (Liu et al., 2024a) adopts GPT-3.5 (Brown et al., 2020) for the summarizing task. Despite their notable success to capture the richer item semantics, they face a trade-off between efficiency and effectiveness when representing items in natural language, leading us to utilize image information to address this trade-off. Meanwhile, VIP5 (Geng et al., 2023) and UniMP (Wei et al., 2024) have attempted to incorporate image information into LLMs. However, they lack an in-depth investigation into the overlap between image-description pairs and fail to provide effective alignment strategies between image and language spaces tailored to the recommendation context.
6. Conclusion
In this work, we address the trade-off between efficiency and effectiveness when representing item semantics in natural language for an LLM-based recommender system. Based on the observation that there exists a significant information overlap between images and descriptions associated with items, we propose I-LLMRec, a novel method that leverages images to capture rich item semantics of lengthy descriptions with only a few tokens, thereby capturing user preferences efficiently and effectively. However, as the image and language space are not inherently aligned, we propose the Image-LLM Alignment (ILA) mode, which effectively bridges the gap between these spaces. Furthermore, we propose the Image-based REtrieval (IRE) mode, a retrieval-based recommendation approach, to enhance both the reliability and efficiency of the recommender systems. Our extensive experiments demonstrate that I-LLMRec outperforms natural language-based approaches in terms of both efficiency and effectiveness.
Appendix A Ethics Statement
To the best of our knowledge, we comply with the KDD Code of Ethics and do not violate any ethical issues. Throughout this paper, we use publicly available data and code.
Figure 8. Prompt for summarizing descriptions. The red-colored text denotes the actual metadata for the item.
Appendix B Prompt for Summarization
Figure 8 shows the prompt used for description summarization. Following TRSR (Zheng et al., 2024), we incorporate attribute information along with the description to generate a concise summary.
Appendix C Recommendation Protocol
In this section, we describe a comprehensive explanation of the recommendation protocol, which is consistently applied to Attribute-based Representation, Description-based Representation (i.e., Full Description and Summarized Description), and I-LLMRec as introduced in Section 1. Note that this recommendation protocol follows a retrieval-based recommendation approach similar to the IRE module discussed in Section 3.3. Therefore, we recommend reading Section 3.3 first for a deeper understanding. At a high level, we modify the way items are represented by replacing the visual representation with either an attribute-based or description-based representation within Equation 2, followed by a recommending process similar to IRE module.
[…]
Figure 9. Cosine similarity between item image-description pairs in Amazon Phone and Grocery datasets and image-caption pairs in the COCO dataset using CLIP (Radford et al., 2021).
Figure 10. Sigmoid-based similarity (ranging from 0 to 1) between item image-description pairs across four datasets (Amazon Sport, Art, Grocery, and Phone datasets) using SigLIP (Zhai et al., 2023).
Appendix D Consistent Observation of Information Overlap
We further investigate the information overlap between image and description associated with items, which was explored in Section 1, by expanding these experiments to additional datasets and leveraging another CLIP=based model. 1) Expanding to additional datasets: Building upon our previous comparative experiments in Amazon Sport and Art—where we used CLIP (Radford et al., 2021) to measure cosine similarity in Section 1—we extend our analysis to the Amazon Grocery and Phone datasets. As shown in Figure 9, the average similarity for Amazon Phone and Grocery is approximately 0.31, which is higher than 0.26 average similarity observed in well-aligned COCO image-caption pairs. This consistent observation across multiple datasets further supports the information overlap between item images and descriptions. 2) Leveraging another CLIP model: To further explore the information overlap using a different CLIP variant, we select SigLIP (Zhai et al., 2023), a variant of CLIP with sigmoid loss function, to compute the similarity scores ranging from 0 to 1. Specifically, when we apply SigLIP to measure similarity between item image-description pairs across all four datasets (Amazon Sport, Art, Grocery, and Phone), we observed that, as show in Figure 10, the majority of items exhibit similarity scores close to 1, further demonstrating the significant information overlap between their image and description pairs.
Table 4. Performance comparison when LLM backbone of  I-LLMRec is fine-tuned using LoRA (Hu et al., 2021).
Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
LLM Frozen 0.4570 0.3711 0.5883 0.4839
LLM fine-tuning with LoRA 0.4633 0.3772 0.5958 0.4949
Appendix E Frozen LLM vs LoRA Fine-Tuned LLM
Note that fine-tuning the LLM backbone is not mandatory for I-LLMRec. Nevertheless, in this section, we explore potential performance gains from fine-tuning the LLM, a common approach in existing LLM-based recommender systems (Lin et al., 2024; Bao et al., 2023; Tan et al., 2024; Zheng et al., 2024; Liu et al., 2024a). However, since fully fine-tuning an LLM demands substantial computational resources, we adopt the parameter-efficient fine-tuning strategy, LoRA (Hu et al., 2021), which introduces a small number of learnable rank matrices into each LLM’s transformer layer. As shown in Table 4, we observed the performance improvement through fine-tuning, indicating that there is room for further improvement within I-LLMRec. However, considering the computational cost, we opt to freeze the LLM to preserve its intrinsic knowledge while improving training efficiency.
Table 5. Data statistics after preprocessing.
Dataset # Users # Items # Interactions Avg. Len
Sport 24,665 9,884 179,491 7.28
Grocery 84,896 25,632 739,628 8.71
Art 14,986 5,801 121,917 8.13
Phone 59,307 19,671 403,194 6.80
Appendix F Data Statistics
Table 5 shows the summary of the data statistics after preprocessing. Our experiments were conducted on datasets of varying scale, ranging from Grocery, which has a large number of interactions, to Art, which has relatively fewer interactions.
Table 6. Impact of the different backbones.
LLM Vision Encoder Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
LLaMA-2.7B SigLIP 0.4570 0.3711 0.5883 0.4839
LLaMA-2.7B CLIP 0.4476 0.3560 0.5836 0.4796
Vicuna 1.5-7B SigLIP 0.4670 0.3779 0.5943 0.4930
Vicuna 1.5-7B CLIP 0.4552 0.3677 0.5885 0.4892
Appendix G Impact of Different LLMs and Vision Encoders
To investigate the impact of the backbone models on the recommendation performance of  I-LLMRec, we evaluated our model by replacing different backbone models. For the LLM, we used Sheared LLaMA-2.7B (Xia et al., 2023) and Vicuna 1.5-7B (Chiang et al., 2023), an improved version of LLaMA-7B (Touvron et al., 2023a). For the vision encoder, we used CLIP [^11] https://huggingface.co/openai/clip-vit-large-patch14-336 (Radford et al., 2021) and SigLIP [^12] https://huggingface.co/google/siglip-so400m-patch14-384 (Zhai et al., 2023), with the advanced version of CLIP scaling with sigmoid loss. Table 6 shows the following observations: 1) Given the same vision encoder, I-LLMRec with Vicuna 1.5-7B outperforms the model with LLaMA-2.7B. This improvement is likely attributed to Vicuna 1.5-7B’s superior semantic reasoning capability, allowing it to capture user preferences more effectively. 2) Given the same LLM backbone, I-LLMRec with SigLIP performs better than the version with CLIP. This suggests that models that better capture visual features are more effective in delivering item semantics, thereby enabling the LLM to better understand the item semantics. In summary, employing more powerful LLMs and vision encoders leads to improvements in recommendation performance, as these models effectively capture item semantics and user preference.
Appendix H Training Strategy
Throughout this paper, we optimize our model using an end-to-end training strategy to ensure ease of implementation and learning efficiency. However, in this section, we explore the impact of an alternative two-stage training strategy, where we separate the ILA and IRE modules. Specifically, for stage 1, we train the adaptor $M$ through the ILA module to initially align the visual space with the language space. For stage 2, we freeze the adaptor and train only $F_{u}^{*}$ and $F_{i}^{*}$ (*=[Img, CF, Text]) through the IRE module. As shown in Table 7, the two-stage strategy shows performance comparable to the end-to-end strategy, indicating that neither strategy offers a clear advantage over the other. Given this, we opt for the end-to-end training strategy to facilitate the ease of implementation and training efficiency.
Table 7. Training Strategy.
Training Strategy Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
End-to-End 0.4570 0.3711 0.5883 0.4839
Stage1+Stage2 0.4604 0.3747 0.5858 0.4887
Figure 11. Prompt template for ILA module. The red-colored text denotes the actual information of an item.
Appendix I ILA Prompt Templates
Figure 11 shows the prompt template for ILA module of  I-LLMRec. As described in Section 3.2, we utilize the prompt templates to align the visual feature with the language space of LLMs by training the adaptor network $M$.